# Multi-objective Bayesian optimization for Likelihood-Free inference in sequential sampling models of decision making

### v2

@@ Abstract
Joint modeling of different data sources in decision-making processes is crucial for understanding decision dynamics in consumer behavior models. Sequential Sampling Models (SSMs), grounded in neuro-cognitive principles, provide a systematic approach to combining information from multi-source data, such as those based on response times and choice outcomes. However, parameter estimation of SSMs is challenging due to the complexity of joint likelihood functions. Likelihood-Free inference (LFI) approaches enable Bayesian inference in complex models with intractable likelihoods, like SSMs, and only require the ability to simulate synthetic data from the model. Extending a popular approach to simulation efficient LFI for single-source data, we propose Multi-objective Bayesian Optimization for Likelihood-Free Inference (MOBOLFI) to estimate the parameters of SSMs calibrated using multi-source data. MOBOLFI models a multi-dimensional discrepancy between observed and simulated data, using a discrepancy for each data source. Multi-objective Bayesian Optimization is then used to ensure simulation efficient approximation of the SSM likelihood. The use of a multivariate discrepancy allows for approximations to individual data source likelihoods in addition to the joint likelihood, enabling both the detection of conflicting information and a deeper understanding of the importance of different data sources in estimating individual SSM parameters. We illustrate the advantages of our approach in comparison with the use of a single discrepancy in a simple synthetic data example and an SSM example with real-world data assessing preferences of ride-hailing drivers in Singapore to rent electric vehicles. Although we focus on applications to SSMs, our approach applies to the Likelihood-Free calibration of other models using multi-source data.
@@
, and
@@ 1 Introduction
Understanding individual-level consumer preferences is the basis of estimating and forecasting system-level demand in many applications in marketing and economics. Most individual-level behavioural models assume that the decision-makers are rational and choose an alternative with the highest utility. These models assume that decision-makers can process all available information and map it to the utility of each option before making a utility-maximizing choice. In the absence of data on decision-making processes, validating the assumptions of full information processing and utility-maximising behaviour is infeasible. For instance, these assumptions may not hold when decision-makers face realistic choice situations under time pressure, as they might not process all available information and deviate from utility-maximising decision rules (Caplin, 2016; Clithero, 2018).
With the advent of computer-based surveys and technological innovations, several types of process datasets can be collected, such as response time (i.e., the time elapsed from the presentation of alternative information to the confirmation of choice), mouse motion, and eye movements of the decision-maker (Bansal, Kim and Ozdemir, 2024). Some studies have explored the relationships between process datasets and choice outcomes. For instance, there is evidence of a speed-accuracy trade-off, where choice accuracy is inversely related to response time (Heitz, 2014), and a gaze-cascade effect, where decision-makers spend more time gazing at the chosen alternative before making a decision (Shimojo et al., 2003). These observations provide initial evidence that multi-source data on the decision-making process can enhance understanding of human decisions and offer better explanations of consumer preferences.
However, modelling multi-source data is challenging using the classic utility-based choice models as they are inherently static and do not have mechanisms to model the dynamic aspects of the decision-making process. Two potential approaches exist to model the decision-making process data with choice outcome data. First, a two-step choice model can be developed, such that the information searching process can be modelled using the decision-making process data in the first step, and choice outcome data can be used in the second step to explain the final decision conditional on the information processing (Gilbride and Allenby, 2004; Cheremukhin, Popova and Tutino, 2015). Second, Sequential Sampling Models (SSMs) from cognitive science provide a framework to utilize multi-source data to endogenously model the decision-making process and choice outcome (Smith, 2000). SSMs posit that the decision-maker gradually accumulates noisy information about available alternatives until the evidence reaches a certain threshold. The product with the highest positive evidence is chosen by the decision-maker. Neuroscientists also agree that the brain signals of decision-makers are compatible with the assumed decision-making process in SSMs (Pisauro et al., 2017). Based on the different structural formulations of the evidence accumulation process, several variants of SSMs have emerged, such as Multi-attribute Linear Ballistic Accumulator (MLBA, Trueblood, Brown and Heathcote, 2014) and Multi-alternative decision by sampling Model (MDbS, Noguchi and Stewart, 2018). While the two-step choice model is data-dependent and ad hoc, SSMs provide a generalizable framework to jointly model final choice (CH), response time (RT), and other process datasets.
Even though the framework and applications for SSMs are promising, their bottom-up design generally causes intractable joint likelihood of multi-source data, resulting in challenges associated with parameter estimation. As a compromise strategy, some researchers discretize RT into a few time steps to reduce computational burden (Noguchi and Stewart, 2018). Another strand of studies marginalizes the RT to only fit choice outcome data if the resulting likelihood has a closed-form expression (Ratcliff et al., 2016; Hancock, Hess and Choudhury, 2018), losing the core value of SSMs.
If computational issues are resolved, SSMs have the potential to become a workhorse model in marketing and economics to leverage decision-making process data in eliciting consumer preferences. In the absence of a closed-form likelihood for multi-source data, there is a need to develop a Likelihood-Free inference (LFI) approach to estimate SSMs that is computationally efficient and facilitates checks of consistency between different types of decision processes and choice data. LFI methods perform Bayesian inference using simulation of the model as a replacement for intractable likelihood evaluations. One of the most well-established LFI methods used in SSMs is approximate Bayesian computation (ABC, Sisson, Fan and Beaumont, 2018), which approximates the posterior by parameter samples for which a simulated dataset has a discrepancy from the observed dataset below a certain tolerance. See Section 2.1 for a more detailed discussion. ABC implicitly uses a kernel density approximation to the likelihood for data summaries. A competitive approach, probability density approximation (PDA, Turner and Sederberg, 2014), has also been widely used to estimate SSMs for over a decade. However, both ABC and PDA suffer from the curse of dimensionality, making them computationally intractable when the number of data summary statistics is large, which is often required for highly parameterized models. Structured density estimation approaches which deal better with high-dimensional summary statistics include synthetic likelihood, which is based on a working normal model for summary statistics (Wood, 2010; Frazier et al., 2022), flexible extensions of synthetic likelihood based on copulas or empirical saddle point approximation (Fasiolo et al., 2018; An, Nott and Drovandi, 2020) and neural density estimation approaches (Papamakarios and Murray, 2016; Papamakarios, Sterratt and Murray, 2019; Greenberg, Nonnenmacher and Macke, 2019). Modern neural LFI methods have been used in the context of SSMs (e.g. Radev et al. (2023)) and for multi-source data (Schmitt, Radev and Bürkner, 2023). However, these structured density estimation methods require a large number of model simulations to obtain adequate approximations, which may be infeasible if simulation from the model is computationally expensive. A preferred approach in such settings is to use active learning approaches such as Bayesian optimization for Likelihood-Free inference (BOLFI, Gutmann and Corander, 2016), and extending these methods is the focus of our work. We review the BOLFI approach in Section 2.3, and related existing literature in Section A.2 of the Supplement. Our novel multi-objective extension of BOLFI (hereafter, MOBOLFI) shares the simulation efficiency of BOLFI but has key advantages compared to other Likelihood-Free approaches in handling intractable models for multi-source data. Crucially, our approach has the capability to approximate both the joint likelihood as well as likelihoods for individual data sources, enabling checks of the consistency of information supplied by different parts of the data. Examining this consistency is essential for trustworthy analysis of complex multi-source data such as those arising for SSMs. It is also valuable for understanding what data sources are most informative for the estimation of individual SSM parameters.
@@ 2.1 Approximate Bayesian computation
ABC methods are based on an approximation of the likelihood in terms of a synthetic dataset being close to the observed data according to some discrepancy. The discrepancy is often defined in terms of data summary statistics. It is important to explain the ABC likelihood approximation, since this motivates the BOLFI approach discussed in subsection 2.3. Interested readers will find a more extensive discussion of ABC than we give here in Sisson, Fan and Beaumont (2018).
@@ 2.3 BOLFI
pt(Sobs|θ)
Φ(tμnf(θ)σnf2(θ)+σ2)
@@
[…]
where
[…]
denotes “approximately proportional to”. This likelihood approximation can be used with MCMC sampling to draw approximate posterior samples from $p_{t}(\theta|S_{\text{obs}})$. The choice of tolerance $t$ in our examples is discussed later.
@@ 3.3 MOBOLFI
Gutmann and Corander (2016) chose the tolerance $t$ in the univariate BOLFI method as the $q$-quantile of $\Delta_{1},...,\Delta_{n_{f}}$, where $q\in(0,1)$. In the bivariate MOBOLFI method, we extend the choice of tolerance $t=(t_{1},t_{2})$ to the 2-dimensional vector $q$-quantile of $\Delta_{1},...,\Delta_{n_{f}}$, where $q\in(0,1)^{2}$. The approximate likelihood (3.6) is sensitive to the value of $t$. For results in this paper, we set the $q=0.05$. A comparison study over different $q$-quantile tolerance is given in for 3 different examples in the Supplement. The dependent noise covariance matrix $\Sigma$ is estimated by the covariance of a repeatedly simulated sample $\{\Delta_{i_{\Sigma},j}\}_{j=1}^{n_{\Sigma}}$, where $\Delta_{i_{\Sigma},j}=D(\theta_{i_{\Sigma}})+\epsilon_{i_{\Sigma}}$ are simulated noisy discrepancies for some $\theta_{i_{\Sigma}}$. For results in this paper, we set $n_{\Sigma}=100$ and $\theta_{i_{\Sigma}}=\underset{(\theta_{i},\Delta_{i})\in T_{n_{f}}}{\operatorname*{{arg\,min}}}(\Delta_{i}-\mu_{n_{f}}(\theta_{i})^{\top}\Sigma_{n_{f}}(\theta_{i})^{-1}(\Delta_{i}-\mu_{n_{f}}(\theta_{i}))$. For better performance and numerical stability, we apply scaling to control the magnitude of $\Delta$. Details of scaling are provided in Section A.3 of the Supplement.
@@ 4.2.2 Synthetic data
Figure  2 shows approximate posterior distributions obtained using the BOLFI and MOBOLFI methods. We make a number of observations. First, the variance of the MOBOLFI approximate posterior is smaller than that for BOLFI. In fact, MOBOLFI using only the response time data produces more accurate approximations of the true posterior than BOLFI given both data sources. This suggests that when it is natural to define discrepancies separately for different data sources, combining them linearly, even with weights, may result in information loss. Second, both MOBOLFI and BOLFI posterior variances are larger than for the true posterior, which is expected, due to the irreducible information loss from approximation. Third, when comparing the MOBOLFI posteriors calculated by likelihoods of different data sources, the posterior using both data sources is better than any posterior using only one data source. Fourth, for inference of $\beta$ and $\lambda_{1}$, the posterior using only $CH$ has its marginal maximum a posteriori probability (MAP) value far from the value of $\theta^{\text{true}}$; for inference of $\chi$ the posterior using only $CH$ has variance larger than that of the posterior using the joint likelihood, compared to using $RT$ only or the joint likelihood. In psychology and economics, researchers usually focus on the choice data $CH$ in parameters inference of SSMs like MLBA. The figure shows that using response time data in inference not only reduces the approximate posterior variance to make it closer to the true posterior variance, but also helps locate the area of the global maximum. Such conclusions are likely to extend to other variants of MLBA and LBA. Section C.3 of the Supplement details additional experiments on the factors affecting the performance of MOBOLFI in this example.
@@ 4.2.3 Real-world data
MOBOLFI is applied to infer parameters of MLBA on a real-world dataset from a consumer choice experiment. This experiment assesses the rental preferences of ride-hailing drivers in Singapore for electric vehicles (EVs) via a street-intercept survey (Ding et al., 2024). Before the stated preference experiment, the driver’s basic information such as working days per week and information about currently rented internal combustion engine vehicles (ICEVs) is collected. The drivers were asked to make a choice among three alternatives including ICEVs, Electric Vehicle Model A (EVA), and Electric Vehicle Model B (EVB) with three listed attributes, which are monthly rental cost (RC, in SGD), daily operating cost (OC, in SGD), and daily mileage (DR, in km). The time that elapses from the appearance of information to the confirmation of choice is recorded as RT. Note that EVA and EVB are assumed to be identical except for their values of three listed attributes. After data preprocessing, 149 participants with 584 valid observations are used in the parameter estimation. Since monthly rental cost (RC) and daily operating cost (OC) capture the monetary aspect, these two can be merged into one alternative specific attribute for the monthly total cost (TC, in SGD).
@@
The unknown paramter is $\theta=(\lambda_{1},\log{\beta_{TC}},\log{\beta_{\log{DR}}},\zeta_{ICEV},\log{(\chi-\mathcal{A})})$, and the rest of the parameters are fixed at known values. Firstly, we set $\mathcal{A}=45,s=1,\zeta_{EVB}=0$, and $I_{0}=1$ to ensure parameter identification. Secondly, the alternative-specific constant $\zeta$ is often interpreted as the initial preference of the corresponding alternative before considering the attribute values in empirical studies. A mild and reasonable assumption made here is that $\zeta_{EVA}=\zeta_{EVB}=0$, since both are EVs. Furthermore, according to the posterior of $\lambda_{2}$ using the closed-form likelihood when estimable, it is very close to 0, hence, $\lambda_{2}=0$ is used for the final empirical analysis. Section C.4.1 in the Supplement gives detailed results obtained using the closed-form likelihood. Finally, similarly to synthetic data estimation, $\beta_{TC}$,$\beta_{\log{DR}}$, and $\chi-\mathcal{A}$ are log-transformed to increase sampling efficiency.
@@
Analogous to the synthetic data, for the implementation of MOBOLFI, 100 initial prior samples $\{\theta^{(i)},(RT^{(i)},CH^{(i)})\}_{i=1}^{100}$ are drawn, where $\lambda_{1}^{(i)}\sim\mathcal{U}[0,5]$, $\zeta_{ICEV}^{(i)}\sim\mathcal{U}[-3,5]$, $\log{\beta_{TC}^{(i)}}\sim\mathcal{U}[-3,5]$, $\log{\beta_{\log{DR}}^{(i)}}\sim\mathcal{U}[-3,5]$, and $\log{(\chi-\mathcal{A})}^{(i)}\sim\mathcal{U}[-2,6]$. The training data is $\{\theta^{(i)},(\Delta^{*}_{1}(RT^{(i)},RT^{o}),\Delta^{*}_{2}(CH^{(i)},CH^{o}))\}_{i=1}^{100}$, where $(RT^{o},CH^{o})$ are observed empirical data. Unlike the synthetic data experiment using single simulated RT and CH data for discrepancies by eq. (4.4) and eq. (4.5), $(RT^{(i)},CH^{(i)})=\{(RT^{(i)}_{s},CH^{(i)}_{s})\}_{s=1}^{S}$ is a collection of multiple simulations independently generated by the MLBA simulator with $\theta^{(i)}$ and attribute matrix $X$. The discrepancy for the RT data source $\Delta_{1}^{*}(RT^{(i)},RT^{o})$ is the average of $1$-Wasserstein distance between the empirical distribution of log-transformed $RT^{o}$ and the replicates of $RT^{(i)}$:
$\Delta_{1}^{*}(RT^{(i)},RT^{o})=\frac{\sum_{s=1}^{S}\|\log(\widetilde{RT}^{o}-\log(\widetilde{RT}^{i}_{s}))\|_{1}}{S}.$ (4.7)
@@
The discrepancy $\Delta_{2}^{*}(CH^{(i)},CH^{o})$ for the choice data is the Brier score (Brier, 1950) between the empirical probabilities over the $S=30$ replicates for the synthetic choice data and the observed choice outcomes (i.e. the average squared difference between empirical probabilities by simulations and observed outcomes):
$\Delta^{*}_{2}(CH^{(i)},CH^{o})=\frac{1}{3\times 584}\sum_{j=1}^{584}\bigg{(}CH^{o}_{j}-\frac{\sum_{s=1}^{S}CH^{(i)}_{sj}}{S}\bigg{)}^{2}.$ (4.8)
There are two reasons for adopting $\Delta^{*}$ instead of $\Delta$ in eq.(4.4) and eq.(4.5) for the empirical experiment. Firstly, for the CH discrepancy, the Brier score in eq.(4.8) offers a more informative discrepancy by capturing the empirical choice distribution $\frac{\sum_{s=1}^{S}CH^{(i)}_{s}}{S}$ rather than a chosen instance. Secondly, for both RT and CH discrepancies, $\Delta_{1}^{*}$ and $\Delta_{2}^{*}$ offer more robust discrepancies than $\Delta$ thanks to averaging over $S$ replicates. Consequently, $\Delta_{1}^{*}$ and $\Delta_{2}^{*}$ result in a quicker convergence during MOBOLFI GP model training while requiring S times as many model simulations per acquisition as $\Delta_{1}$ and $\Delta_{2}$. For a comparison between $\Delta$ and $\Delta^{*}$, see Section C.4.2 of the Supplement. De-MCMC is used for both the MOBOLFI approximate posterior and closed-form MLBA posterior estimation.
Figure 5 shows approximate marginal MOBOLFI posteriors for selected parameters and the corresponding posteriors for the closed-form MLBA likelihood. The figure demonstrates that the contributions of different data sources to the estimation of individual parameters can vary greatly according to the parameter, and MOBOLFI can reveal these relationships. For example, the MOBOLFI marginal posterior of $\delta_{ICEV}$ is very diffuse when conditioning on only response-time data, showing that the MOBOLFI posterior conditioning on both data sources is informative mostly because of the choice data. This aligns with the interpretation from psychology that $\delta$ stands for the initially rooted preference for an alternative before taking a survey which is independent of the decision-making process, hence rarely captured by RT data only. On the other hand, the $\log{(\chi-\mathcal{A})}$ MOBOLFI marginal posterior conditioning on both choice and response time is more similar to its marginal posterior conditional on response time only. This is consistent with the interpretation that $\chi$ reflects the decision difficulty and respondent’s cautiousness, which are associated with RT. In addition, the marginal MOBOLFI posteriors of $\log{\beta_{TC}},\log{\beta_{\log{DR}}}$, and $\lambda_{1}$ display large variance and different patterns from the closed-form posterior. Section C.4.3 in the Supplement shows a similar distribution of the discrepancies using the closed-form and joint MOBOLFI MAP estimates, and also demonstrate that the distribution of the discrepancy depends weakly on $\log{\beta_{TC}},\log{\beta_{\log{DR}}}$ and $\lambda_{1}$.

### v3

@@ Abstract
Scientifically motivated statistical models can sometimes be defined by a generative process for simulating synthetic data. Models specified this way can have likelihoods which are intractable, and this is the case for many sequential sampling models (SSMs) widely used in psychology and consumer behavior modelling. Researchers have developed likelihood-free inference (LFI) methods to make Bayesian inferences on parameters in models with intractable likelihood. Extending a popular approach to simulation efficient LFI for single-source data, we propose Multi-objective Bayesian Optimization for Likelihood-Free Inference (MOBOLFI) to estimate the parameters of SSMs calibrated using multi-source data, such as those based on response times and choice outcomes. MOBOLFI models a multi-dimensional discrepancy between observed and simulated data, using a discrepancy for each data source. Multi-objective Bayesian Optimization is then used to ensure simulation efficient approximation of the SSM likelihood. The use of a multivariate discrepancy allows for approximations to individual data source likelihoods in addition to the joint likelihood, enabling both the detection of conflicting information and a deeper understanding of the importance of different data sources in estimating individual SSM parameters. We illustrate the advantages of our approach in comparison with the use of a single discrepancy in a simple synthetic data example and an SSM example with real-world data assessing preferences of ride-hailing drivers in Singapore to rent electric vehicles. Although we focus on applications to SSMs, our approach applies to the likelihood-free calibration of other models using multi-source data.
@@
, thanks: David Chen and Xinwei Li contributed equally to this work., , and
@@ 1 Introduction
In many scientific applications, the most natural way to incorporate subject matter knowledge into a statistical model is by specifying a generative algorithm (or simulator) for synthetic data. In this case, the likelihood function can sometimes be difficult to compute, and this makes traditional statistical inference methods hard to apply. In response to this challenge, researchers have developed likelihood-free inference (LFI) methods for conducting Bayesian inference using only simulation from the model, without any likelihood evaluations. An alternative name for LFI is simulation-based inference (SBI), which stresses the role of simulating synthetic data.
Our work considers settings where it is important to obtain accurate approximations to the posterior density using as few simulations as possible, and where the likelihood factorizes into terms for different data sources. We extend a widely-used simulation efficient method for single source data, Bayesian optimization for likelihood-free inference (BOLFI) (Gutmann and Corander, 2016). BOLFI uses Bayesian optimization to sequentially choose the simulation parameters to make best use of a limited computational budget. In the case of multi-source data, we consider a multi-objective extension which we call MOBOLFI, which measures fidelity between synthetic and observed data by a multi-dimensional discrepancy with a component for each data source. The multi-dimensional discrepancy is modelled, and this allows approximation of likelihood terms for different data, which is important for understanding the importance of different data sources for inference about individual parameters and for model checking.
There are many existing methods for LFI. One of the most well-established methods is approximate Bayesian computation (ABC, Sisson, Fan and Beaumont, 2018), which approximates the posterior by parameter samples for which a simulated dataset has a discrepancy from the observed dataset below a certain tolerance. See Section 2.1 for a more detailed discussion. ABC implicitly uses a kernel density approximation to the likelihood for data summaries. A competitive approach, probability density approximation (PDA, Turner and Sederberg, 2014), has also been widely used for over a decade. However, with high-dimensional summary statistics, both ABC and PDA can perform poorly, making then unsuitable for many highly parameterized models.
Structured density estimation approaches which deal better with high-dimensional summary statistics include synthetic likelihood, which is based on a working normal model for summary statistics (Wood, 2010; Frazier et al., 2022), flexible extensions of synthetic likelihood based on copulas or empirical saddle point approximation (Fasiolo et al., 2018; An, Nott and Drovandi, 2020) and neural density estimation approaches (Papamakarios and Murray, 2016; Papamakarios, Sterratt and Murray, 2019; Greenberg, Nonnenmacher and Macke, 2019). However, these structured density estimation methods require a large number of model simulations to obtain adequate approximations, which may be infeasible if simulation from the model is computationally expensive. A preferred approach in such settings is to use active learning approaches such as Bayesian optimization for Likelihood-Free inference (BOLFI, Gutmann and Corander, 2016), and extending these methods is the focus of our work. We review the BOLFI approach in Section 2.3, and related existing literature in Section A.2 of the Supplement.
Another challenge for LFI methods is model misspecification (Wilkinson, 2013; Frazier, Robert and Rousseau, 2020; Frazier and Drovandi, 2021; Ward et al., 2022). There is an extensive literature on this issue, but here we focus on model checking with multi-source data, where the different data sources may bring conflicting information about the model parameter (Presanis et al., 2013). Modern LFI methods have been used in the context of Sequential Sampling Models (e.g. Radev et al. (2023)) and for multi-source data (Schmitt, Radev and Bürkner, 2023; Bahg et al., 2020), but questions of model adequacy have been previously addressed in a global way. The MOBOLFI method proposed here has the capability to approximate both the joint likelihood as well as likelihoods for individual data sources, enabling checks of the consistency of information supplied by different parts of the data. Examining this consistency is essential for trustworthy analysis of complex multi-source data such as those arising for Sequential Sampling Models (SSMs). It is also valuable for understanding which data sources are most informative for the estimation of individual SSM parameters.
@@ 2.1 Approximate Bayesian computation
ABC methods approximate the likelihood by the probability of a synthetic dataset being close to the observed data in terms of some discrepancy. The discrepancy is often defined in terms of data summary statistics. It is important to explain the ABC likelihood approximation, since this motivates the BOLFI approach discussed in subsection 2.3. For a more extensive discussion of ABC than we give here see Sisson, Fan and Beaumont (2018).
@@ 2.3 BOLFI
$p_{t}(S_{\text{obs}}|\theta)\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}\Phi\left(\frac{t-\mu_{n_{f}}(\theta)}{\sqrt{\sigma_{n_{f}}^{2}(\theta)+\sigma^{2}}}\right),$
@@ 3.3 MOBOLFI
Gutmann and Corander (2016) chose the tolerance $t$ in the univariate BOLFI method as the $q$-quantile of $\Delta_{1},...,\Delta_{n_{f}}$, where $q\in(0,1)$. In the bivariate MOBOLFI method, we extend the choice of tolerance $t=(t_{1},t_{2})$ to the 2-dimensional vector $q$-quantile of $\Delta_{1},...,\Delta_{n_{f}}$, where $q\in(0,1)^{2}$. The approximate likelihood (3.6) is sensitive to the value of $t$. For results in this paper, we set the $q=0.05$. A comparison study over different $q$-quantile tolerance is given for 3 different examples in the Supplement. The dependent noise covariance matrix $\Sigma$ is estimated by the covariance of a repeatedly simulated sample $\{\Delta_{i_{\Sigma},j}\}_{j=1}^{n_{\Sigma}}$, where $\Delta_{i_{\Sigma},j}=D(\theta_{i_{\Sigma}})+\epsilon_{i_{\Sigma}}$ are simulated noisy discrepancies for some $\theta_{i_{\Sigma}}$. For results in this paper, we set $n_{\Sigma}=100$ and $\theta_{i_{\Sigma}}=\underset{(\theta_{i},\Delta_{i})\in T_{n_{f}}}{\operatorname*{{arg\,min}}}(\Delta_{i}-\mu_{n_{f}}(\theta_{i})^{\top}\Sigma_{n_{f}}(\theta_{i})^{-1}(\Delta_{i}-\mu_{n_{f}}(\theta_{i}))$. For better performance and numerical stability, we apply scaling to control the magnitude of $\Delta$. Details of scaling are provided in Section A.3 of the Supplement.
@@ 4.2.2 Synthetic data
Figure 2 shows approximate posterior distributions obtained using the BOLFI and MOBOLFI methods. We make a number of observations. First, the variance of the MOBOLFI approximate posterior is smaller than that for BOLFI. In fact, MOBOLFI using only the response time data produces more accurate approximations of the true posterior than BOLFI given both data sources. This suggests that when it is natural to define discrepancies separately for different data sources, combining them linearly, even with weights, may result in information loss. Second, both MOBOLFI and BOLFI posterior variances are larger than for the true posterior, which is expected, due to the irreducible information loss from approximation. Third, when comparing the MOBOLFI posteriors calculated by likelihoods of different data sources, the posterior using both data sources is better than any posterior using only one data source. Fourth, for inference of $\beta$ and $\lambda_{1}$, the posterior using only $CH$ has its marginal maximum a posteriori probability (MAP) value far from the value of $\theta^{\text{true}}$; for inference of $\chi$ the posterior using only $CH$ has variance larger than that of the posterior using the joint likelihood, compared to using $RT$ only or the joint likelihood. In psychology and economics, researchers usually focus on the choice data $CH$ in parameter inference for SSMs like MLBA. The figure shows that using response time data in inference not only reduces the approximate posterior variance to make it closer to the true posterior variance, but also helps locate the area of the global maximum. Such conclusions are likely to extend to other variants of MLBA and LBA. Section C.3 of the Supplement details additional experiments on the factors affecting the performance of MOBOLFI in this example.
@@ 4.2.3 Real-world data
MOBOLFI is applied to infer parameters of MLBA on a real-world dataset from a consumer choice experiment. This experiment assesses the rental preferences of ride-hailing drivers in Singapore for electric vehicles (EVs) via a street-intercept survey (Ding et al., 2024). Before the stated preference experiment, the driver’s basic information such as working days per week, and information about currently rented internal combustion engine vehicles (ICEVs) is collected. The drivers were asked to make a choice among three alternatives including ICEVs, Electric Vehicle Model A (EVA), and Electric Vehicle Model B (EVB) with three listed attributes, which are monthly rental cost (RC, in SGD), daily operating cost (OC, in SGD), and daily mileage (DR, in km). The time that elapses from the appearance of information to the confirmation of choice is recorded as RT. Note that EVA and EVB are assumed to be identical except for their values of three listed attributes. After data preprocessing, 149 participants with 584 valid observations are used in the parameter estimation. Since monthly rental cost (RC) and daily operating cost (OC) capture the monetary aspect, these two can be merged into one alternative specific attribute for the monthly total cost (TC, in SGD).
@@
The unknown parameter is $\theta=(\lambda_{1},\log{\beta_{TC}},\log{\beta_{\log{DR}}},\delta_{ICEV},\log{(\chi-\mathcal{A})})$, and the rest of the parameters are fixed at known values. We set constants $\mathcal{A}=45,s=1,\delta_{EVB}=0$, and $I_{0}=1$ to ensure parameter identification. The alternative-specific constant $\delta$ is often interpreted as the initial preference of the corresponding alternative before considering the attribute values in empirical studies. A mild and reasonable assumption made here is that $\delta_{EVA}=\delta_{EVB}=0$, since both are EVs. Furthermore, based on prospect theory that consumers pay more attention to loss than gain (Camerer et al., 1998), $\lambda_{2}$ is normally smaller than $\lambda_{1}$. Therefore, to simplify the estimation, we set $\lambda_{2}=0$ for the final empirical analysis. Finally, similarly to synthetic data estimation, $\beta_{TC}$, $\beta_{DR}$ and $\chi-\mathcal{A}$ are log-transformed to increase sampling efficiency.
@@
We use independent priors $\lambda_{1}\sim\mathcal{U}[0,5]$, $\delta_{ICEV}\sim\mathcal{U}[-3,5]$, $\log{\beta_{TC}}\sim\mathcal{U}[-3,5]$, $\log{\beta_{DR}}\sim\mathcal{U}[-3,5]$, and $\log{(\chi-\mathcal{A})}\sim\mathcal{U}[-2,6]$. For the implementation of MOBOLFI, 282 initial samples $\{\theta^{(i)},(RT^{(i)},CH^{(i)})\}_{i=1}^{282}$ are obtained using a space-filling design (Santner, Williams and Notz, 2018). The training data is $\{\theta^{(i)},(\Delta^{*}_{1}(RT^{(i)},RT^{o}),\Delta^{*}_{2}(CH^{(i)},CH^{o}))\}_{i=1}^{282}$, where $(RT^{o},CH^{o})$ are observed empirical data. Unlike the synthetic data experiment where a single simulated dataset is used for computing discrepancies at each parameter by eq. (4.4) and eq. (4.5), $(RT^{(i)},CH^{(i)})=\{(RT^{(i)}_{s},CH^{(i)}_{s})\}_{s=1}^{S}$ is a collection of multiple simulations independently generated by the MLBA simulator with $\theta^{(i)}$ and attribute matrix $X$. The discrepancy for the RT data source $\Delta_{1}^{*}(RT^{(i)},RT^{o})$ is the log-average of $1$-Wasserstein distance between the empirical distribution of log-transformed $RT^{o}$ and the $S=50$ replicates of $RT^{(i)}_{s}$:
$\Delta_{1}^{*}(RT^{(i)},RT^{o})=\log\Bigg{(}\frac{\sum_{s=1}^{S}\|\log(\widetilde{RT}^{o})-\log(\widetilde{RT}^{(i)}_{s}))\|_{1}}{3S}\Bigg{)},$ (4.7)
where $\widetilde{RT}$ is the vector of order statistics of $RT$.
@@
The discrepancy $\Delta_{2}^{*}(CH^{(i)},CH^{o})$ for the choice data is the log-average of SSE of choice proportion over the $S$ replicates for the synthetic choice data and the observed choice outcomes (i.e. the average squared difference between empirical choice proportions by simulations and observed outcomes):
$\Delta^{*}_{2}(CH^{(i)},CH^{o})=\log\Bigg{(}\frac{\sum_{s=1}^{S}\mathds{1}_{N}^{T}(CH_{s}^{(i)}-CH^{o})(CH_{s}^{(i)}-CH^{o})^{T}\mathds{1}_{N}}{9SN^{2}}\Bigg{)},$ (4.8)
where $N=584$ is the number of observations and $\mathds{1}_{N}\in\mathbb{R}^{N\times 1}$ is the vector whose each element is 1.
There are two reasons for adopting $\Delta^{*}$ instead of $\Delta$ in eq.(4.4) and eq.(4.5) for the empirical experiment. First, log transformation on both discrepancies encourages the MOBOLFI to explore more in the high-density area (when $\Delta$ is closed to 0), and therefore, letting the approximate log-likelihood be more sensitive (larger gradient) around the high-density area in parameter space. Second, the simulation data from $\mathcal{S}$ replicates provide a more robust discrepancy for each data source by reducing randomness brought by the MLBA simulator. De-MCMC is used for both the MOBOLFI approximate posterior and closed-form MLBA posterior estimation.